OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (anonymous):

2x+y=1 is like that right if it is i can help 4x+2y=-2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a.x = 0 b.x = 2 c.There is no x value as there is no solution. d.x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

wait okay i will answer it :)

OpenStudy (anonymous):

its C right?

OpenStudy (anonymous):

2x+y=1 4x+2(-2x+1)=-2 -2x -2x 4x-4x+2=-2 answer to solve y=-2x +1 that is Y -2 -2 the x is 0 =-4

OpenStudy (anonymous):

C